# GraphAny: A Foundation Model for Node Classification on Any Graph

### v2

@@ Abstract
Foundation models that can perform inference on any new task without requiring specific training have revolutionized machine learning in vision and language applications. However, applications involving graph-structured data remain a tough nut for foundation models, due to challenges in the unique feature- and label spaces associated with each graph. Traditional graph ML models such as graph neural networks (GNNs) trained on graphs cannot perform inference on a new graph with feature and label spaces different from the training ones. Furthermore, existing models learn functions specific to the training graph and cannot generalize to new graphs. In this work, we tackle these two challenges with a new foundational architecture for inductive node classification named GraphAny. GraphAny models inference on a new graph as an analytical solution to a LinearGNN, thereby solving the first challenge. To solve the second challenge, we learn attention scores for each node to fuse the predictions of multiple LinearGNNs. Specifically, the attention module is carefully parameterized as a function of the entropy-normalized distance-features between multiple LinearGNNs predictions to ensure generalization to new graphs. Empirically, GraphAny trained on the Wisconsin dataset with only 120 labeled nodes can effectively generalize to 30 new graphs with an average accuracy of 67.26% in an inductive manner, surpassing GCN and GAT trained in the supervised regime, as well as other inductive baselines.
@@ 1 Introduction
Figure 1: Average performance on 31 datasets. GraphAny is trained on a single dataset (Wisconsin or Arxiv) and performs inductive inference on any graph. The other transductive methods need to be trained on each dataset.
Foundation models [1, 39, 40] pre-trained on massive data have drastically changed the landscape of artificial intelligence by their ability to solve any new task conditioned on just a few demonstrations [8, 3]. So far, this ability has mainly been limited to modalities like text or images. An intrinsic property of these modalities is the presence of a shared input space across all tasks (e.g. a vocabulary of tokens or a patch of pixels), providing a natural basis for learning a single foundation model over a broad class of different tasks and generalizing to new ones. Graphs, however, do not possess such a property. Taking the node classification task as an example, each graph may have unique dimensions and semantics for its feature and label spaces (e.g. continuous or discrete), which prevents us from developing graph foundation models [26] in the same way as for the above modalities.
Recently, several attempts have been made to leverage large language models (LLMs) to solve graph tasks [50, 38, 9]. While these methods demonstrate different ways of incorporating graphs into existing LLMs, they can only deal with text-attributed graphs, a subset of graphs whose node features and labels can be easily verbalized as textual sequences and fed into an LLM. For general graphs, node features and labels often do not have textual descriptions, and features may even be continuous, which often renders LLM-based approaches inadequate as graph foundational models.
@@
Therefore, we are interested in an architecture that can solve node classification on any new graph with any feature and label spaces in a “native” manner. Such a goal is challenging for existing models on graphs due to two reasons: (1) Existing models learn transformations specific to the dimension, type, and structure of the features and labels used in the training, and cannot perform inference on feature and label spaces that are different from the training ones. This requires us to develop a new model architecture for arbitrary feature and label spaces. (2) Existing models learn functions specific to the training graph and cannot generalize to new graphs; this calls for an inductive function that can generalize to any graph once it is trained.
Our Contributions
@@
In this work, we present GraphAny, a novel pre-trained foundation model architecture that can solve node classification on any new graph. GraphAny consists of two components: a LinearGNN that performs inference on new feature and label spaces without training steps, and an attention vector for each node based on entropy-normalized distance features that ensure generalization to new graphs. Specifically, our LinearGNN models the mapping between node features and labels as a non-parameteric graph convolution followed by a linear layer, whose parameters are determined in closed form without requiring explicit training steps. While a single LinearGNN model may be far from optimal for many graphs, we employ multiple LinearGNN models with different graph convolution operators and learn an attention vector to fuse their predictions. The attention vector is carefully parameterized as a function of distance features between the predictions of LinearGNNs, which guarantees the model to be invariant to the permutations of feature and label dimensions. To further improve the generalization ability of our model, we propose entropy normalization to rectify the distance feature distribution to a fixed entropy, which reduces the effect of different label dimensions on distance features. Intuitively, the attention vector learns to select the most effective combination of LinearGNNs for each node based on their prediction distributions, reflects statistics of its local structure (e.g., homophily metrics [25]), and generalizes to new graphs.
Empirically, we show that GraphAny trained on any single node classification dataset can generalize to 30 new graphs of different feature and label spaces, even surpassing the average performance of GCN and GAT baselines supervisedly trained on each dataset. We attribute this surprising performance to the attention function GraphAny learns to combine LinearGNNs, which generalizes to new graphs thanks to our careful parameterization. By visualizing the attention in GraphAny, we provide insights into how the attention selects the optimal LinearGNN and how it is affected by the choice of the training graph. We further conduct ablation studies on our entropy normalization and attention parameterization to verify their effectiveness in learning inductive attention. To summarize, the contribution of GraphAny are three folds:
  • We propose LinearGNN, an efficient architecture for inductive node classification on any graph with arbitrary feature and label spaces.
  • We devise an inductive attention module to fuse LinearGNNs using distance features and entropy normalization, which is invariant to feature and label permutations and robust to dimension changes.
  • Combining LinearGNNs and the inductive attention module, GraphAny trained on a single dataset can generalize to any graph, enabling the first graph foundation model for node classification.
@@ 2 Related Work
Inductive Node Classification. It is common to categorize prediction tasks into transductive and inductive setups. In the transductive (semi-supervised) node classification setup, test nodes belong to the same training graph the model was trained on. In the inductive setup, the test graph might be different. The majority of GNNs aiming to work in the inductive setup use known node labels as features. Hang et al. [16] introduced Collective Learning GNNs with iterative prediction refinement. Jang et al. [20] applied a diffusion model for such prediction refinement. Structured Proxy Networks [31] combined GNNs and conditional random fields (CRF) in the structured prediction framework. However, the train-test setup in all those works is limited to different partitions of the same bigger graph, that is, they cannot generalize to unseen graphs with different input feature dimensions and number of classes. To the best of our knowledge, GraphAny is the first approach for inductive node classification to do so.
@@
Labels as Features. Addanki et al. [2] demonstrated that node labels used as features yield noticeable improvements in the transductive node classification setup on MAG-240M in OGB-LSC [18]. Sato [34] used labels as features for training-free transductive node classification with a specific GNN weight initialization strategy mimicking label propagation (hence only applicable to homophilic graphs). In homophilic graphs, node label largely depends on the labels of its close neighbors whereas in heterophilic graphs, node label does not depend on the neighboring nodes. GraphAny supports both homophilic and heterophilic graphs in both transductive and inductive setups.
Graph Foundation Models. Inspired by the progress in language and vision foundation models, the idea of having a single model transferable to an arbitrary graph and, possibly, to different graph-level tasks, has recently been gaining increased attention in the graph learning community. Mao et al. [26] identify the main challenge of graph foundation models (GFM) as finding an invariant graph vocabulary that transfers across different graphs. A variety of methods [50, 38, 9, 11, 29] resort to LLMs as universal featurizers by verbalizing the (sub)graph structure and corresponding task as natural language text. Liu et al. [24] further add a GNN on top of LLM features. However, it is unclear whether the sequential nature of LLMs is suitable for graphs with permutation invariance.
Leveraging GNNs as a backbone of GFMs has been found promising in molecular learning [22, 37], geometric learning for crystal structures [36], and machine learning potentials [4, 23, 48, 45] where a unified graph vocabulary exists, that is, atoms in the periodic table. In non-featurized graphs, GNNs with labeling tricks [49] serve as generalizable link predictors in homogeneous graphs [10] and multi-relational knowledge graphs [13]. GFMs for node classification on featurized graphs present a challenge due to heterogeneous feature and label spaces – GraphAny is the first approach to tackle this problem where a single model can perform node classification on unseen homophilic and heterophilic graphs with various feature types and dimensions and number of labels.
3 GraphAny: A Foundation Model for Node Classification
Figure 2: The goal of graph foundation model (GFM) for node classification: A GFM trained on a pre-training graph $G$ should be able to perform inference on any new graph $G^{\prime}$ with new feature and label spaces.
@@ 3 GraphAny: A Foundation Model for Node Classification
Our goal is to devise a graph foundation model that can perform inductive inference on any new graph with arbitrary feature and label spaces, typically different from the ones associated with the training graph (Figure 2). Here we propose such a solution GraphAny, which consists of two main components (Figure 3): a LinearGNN and an attention module. LinearGNNs provide a basic solution to inductive inference on new graphs with arbitrary feature and label spaces, while the attention module learns to combine multiple LinearGNNs based on inductive features that generalize to new graphs.
@@
In a semi-supervised node classification task, we are given a graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, typically represented as an adjacency matrix ${\bm{A}}$, and node features ${\bm{X}}\in{\mathbb{R}}^{|{\mathcal{V}}|\times d}$, a set of labeled nodes ${\mathcal{V}}_{L}$ and their labels ${\bm{Y}}_{L}\in{\mathbb{R}}^{|{\mathcal{V}}_{L}|\times c}$, where $c$ is the number of unique label classes. The goal of node classification is to predict the labels $\hat{{\bm{Y}}}$ for all the unlabeled nodes ${\mathcal{V}}_{U}={\mathcal{V}}\setminus{\mathcal{V}}_{L}$ in the graph. In the conventional transductive learning setup, this is performed by training a GNN (e.g. GCN [21], GAT [42]) on the subset of labeled nodes using standard backpropagation requiring multiple gradient steps. Such a GNN assumes the full graph to be given and typically does not generalize to a new graph “out of the box” without re-training or at least fine-tuning. Conversely, a graph foundation model is expected to predict the labels $\hat{{\bm{Y}}}$ for any graph without expensive gradient steps. Furthermore, when a new graph is provided, it might have different dimensionality $d^{\prime}$ of the features and number of class labels, $c^{\prime}$, which can also appear in an arbitrary permuted order.
3.1 LinearGNNs for Inductive Inference on Any Graph
@@
A key idea of this paper is to use simple GNN models whose parameters can be expressed analytically. Among existing models, simple graph convolutional networks (SGC) [44] use a non-parameteric graph convolution operation to process node features, followed by a linear layer to predict the labels,
@@
where ${\bm{F}}={\bm{A}}^{k}{\bm{X}}\in{\mathbb{R}}^{|{\mathcal{V}}|\times d}$ is the processed features and ${\bm{W}}\in{\mathbb{R}}^{d\times c}$ is the weight of the linear layer. Originally, the parameters of SGC were trained to minimize a cross-entropy loss on node classification, which does not have analytical solution and requires gradient steps. However, if we assume the label distribution to be Gaussian, minimizing the cross-entropy loss is equivalent to minimizing the mean-squared error loss [14]:
@@
where we use $\hat{{\bm{Y}}_{L}}$ to denote model predictions on the set of labeled nodes. The benefit of this approximation is that now we have a closed-form solution for optimal weights ${\bm{W}}^{*}$:
@@
We term this architecture LinearGNN, as it approximates the prediction of linear GNNs (like SGC) with a single forward pass. Its advantage is that it does not require training and have the same time complexity as a standard GCN at inference time (see Section 3.3). Note that the derivation of LinearGNNs is independent of the graph convolution operation, and can be applied to other linear convolution operations as well. We stress that while we don’t expect LinearGNNs to outperform existing transductive models on node classification, they provide a simple basic module for inductive inference. In Section 3.2, we will see how to combine multiple LinearGNNs to create a stronger model with inductive attention that generalizes to new graphs.
3.2 Learning Inductive Attention over LinearGNNs
LinearGNNs provide a basic solution to inductive inference on new graphs, but they do not learn functions from their training graphs. Besides, our experiments (Figure 6) suggest that different graphs may require LinearGNNs with different convolution operations. Hence, a natural way to incorporate learning in GraphAny is to add an attention module over a set of multiple LinearGNNs. Let $\alpha_{u}^{(1)},\alpha_{u}^{(2)},...,\alpha_{u}^{(t)}$ denote the attention over $t$ LinearGNNs. We generate the final prediction as a combination of all LinearGNN predictions:

### v3

# Fully-inductive Node Classification on Arbitrary Graphs

@@ Abstract
One fundamental challenge in graph machine learning is generalizing to new graphs. Many existing methods following the inductive setup can generalize to test graphs with new structures, but assuming the feature and label spaces remain the same as the training ones. This paper introduces the fully-inductive setup, where models should perform inference on arbitrary test graphs with new structures, feature and label spaces. We propose GraphAny as the first attempt at this challenging setup. GraphAny models inference on a new graph as an analytical solution to a LinearGNN, which can be naturally applied to graphs with any feature and label spaces. To further build a stronger model with learning capacity, we fuse multiple LinearGNN predictions with learned inductive attention scores. Specifically, the attention module is carefully parameterized as a function of the entropy-normalized distance features between pairs of LinearGNN predictions to ensure generalization to new graphs. Empirically, GraphAny trained on a single Wisconsin dataset with only 120 labeled nodes can generalize to 30 new graphs with an average accuracy of 67.26%, surpassing not only all inductive baselines, but also strong transductive methods trained separately on each of the 30 test graphs.
@@ 1 Introduction
Figure 1: Average performance on 31 datasets. GraphAny is trained on a single dataset (Wisconsin or Arxiv) and performs inductive inference on any graph. The other methods have to be trained on each dataset.
One of the most important requirements for machine learning models is the ability to generalize to new data. Models with better generalization abilities are able to perform better on unseen data and tasks, which is a key property for foundation models (Achiam et al., 2023, Team et al., 2023, Touvron et al., 2023) that are designed to accomplish a wide range of downstream tasks. For graphs, generalization is challenging since different graphs usually have different structures and associated attributes. Ideally, graph machine learning models are expected to accommodate this difference and learn functions that are applicable to all graphs.
Many previous works on graphs (Hamilton et al., 2017, Hang et al., 2021, Qu et al., 2022, Jang et al., 2023) consider generalization in the inductive setup, where models are supposed to perform inference on test graphs with new structures different from the training ones. However, these works rely on the assumption that the training and test graphs share the same feature and label spaces, which limits their applications to graphs in a fixed domain (e.g. social networks, citation networks). Ideally, we would like to have a model that generalizes to arbitrary graphs, involving new structures, new dimensions and semantics for their feature and label spaces without additional training. We name this more general and practical setting as the fully-inductive setup (visualized in Figure 2).
Figure 2: Fully-inductive node classification: Trained on a graph $G$, a fully-inductive model should generalize to any new graph $G^{\prime}$ with new feature and label spaces without additional training.
@@
The fully-inductive setup is particularly challenging for existing graph machine learning models for two reasons: (1) Existing models learn transformations specific to the dimension, type, and structure of the features and labels used in the training, and cannot perform inference on feature and label spaces that are different from the training ones. This requires us to develop a new model architecture for arbitrary feature and label spaces. (2) Existing models learn functions specific to the training graph and cannot generalize to new graphs. This calls for an inductive function that can generalize to any graph once it is trained. Despite these challenges, it is always possible to cheat the fully-inductive setup by training a separate instance of existing models for each test dataset. We emphasize that this solution does not solve the fully-inductive setup. However, this cheating solution can be regarded as a strong baseline for fully-inductive models, since it additionally leverages back propagation on the labeled nodes and hyperparameter tuning on the validation data of the test datasets.
@@
Our Contributions. We propose GraphAny to solve node classification in the fully-inductive setup. GraphAny consists of two components: LinearGNNs that perform inference on new feature and label spaces without training steps, and an inductive attention module based on entropy-normalized distance features that ensure generalization to new graphs. Specifically, our LinearGNN models the mapping between node features and labels as a non-parameteric graph convolution followed by a linear layer, whose parameters are determined in analytical form without requiring explicit training steps. While a single LinearGNN model may be far from optimal for many graphs, we employ multiple LinearGNN models with different graph convolution operators and learn an attention vector to adaptively fuse their predictions. The attention vector is carefully parameterized as a function of distance features between the predictions of LinearGNNs, which guarantees the model to be invariant to the permutations of feature and label dimensions. To further improve the generalization ability of our model, we propose entropy normalization to rectify the distance feature distribution to a fixed entropy, which reduces the effect of different label dimensions. By combining both modules, GraphAny learns to combine the LinearGNNs’ predictions for each node based on their prediction distributions, which reflects statistics of its local structure, and generalizes to new graphs.
To summarize, the contribution of GraphAny are four folds:
  • We introduce the fully-inductive setup for generalization across arbitrary graphs. Fully-inductive setup is more general and practical than the conventional inductive setup, allowing knowledge transfer across diverse domains, such as from knowledge graphs to e-commerce graphs.
  • We devise LinearGNN, an analytical solution for node classification that enables efficient inductive inference on any new graph without the need for gradient descent.
  • We identify two necessary properties for fully-inductive generalization: permutation invariance and dimensional robustness w.r.t. the feature and label spaces. Towards these goals, we design an inductive attention module that satisfies these properties and generalizes to new graphs.
  • By combining LinearGNN and the inductive attention module, we present GraphAny, the first fully-inductive model for node classification. Across 31 datasets, GraphAny outperforms strong transductive baselines trained separately on each test dataset and achieves a 2.95$\times$ speedup.
@@ 2 Related Work
Inductive Node Classification. Depending on the test graphs which models generalize to, tasks on graphs can be categorized into transductive and inductive setups. In the transductive (semi-supervised) node classification setup (Kipf and Welling, 2017), test nodes belong to the same training graph the model was trained on. In the inductive setup (Hamilton et al., 2017), the test graph might be different. The majority of GNNs aiming to work in the inductive setup use known node labels as features. Hang et al. (2021) introduced Collective Learning GNNs with iterative prediction refinement. Jang et al. (2023) applied a diffusion model for such prediction refinement. Structured Proxy Networks (Qu et al., 2022) combined GNNs and conditional random fields (CRF) in the structured prediction framework. However, the train-test setup in all those works is limited to different partitions of the same bigger graph, that is, they cannot generalize to unseen graphs with different input feature dimensions and number of classes. To the best of our knowledge, GraphAny is the first approach for fully-inductive node classification, which generalizes to unseen graphs with arbitrary feature and label spaces.
@@
Labels as Features. Addanki et al. (2021) demonstrated that node labels used as features yield noticeable improvements in the transductive node classification setup on MAG-240M in OGB-LSC (Hu et al., 2021a). Sato (2024) used labels as features for training-free transductive node classification with a specific GNN weight initialization strategy mimicking label propagation (hence only applicable to homophilic graphs). In homophilic graphs, node label largely depends on the labels of its close neighbors whereas in heterophilic graphs, node label does not depend on the neighboring nodes. GraphAny supports both homophilic and heterophilic graphs in both transductive and inductive setups.
Connections to Graph Foundation Models. Graph foundation models (GFMs), which aim to develop a single model transferable to new graphs and diverse graph tasks, have garnered significant attention in the graph learning community. The core challenge in designing a GFM is achieving generalization across graphs with varying input and output spaces, which requires identifying an invariant feature space that transfers effectively across graphs (Mao et al., 2024). While fine-tuning a pre-trained model on a new graph is feasible (Zhu et al., 2024), we focus on the more challenging fully-inductive setting, where the model must generalize to unseen graphs without additional training.
In scenarios where graphs share a common feature space, such as in molecular learning (Kläser et al., 2024, Kovács et al., 2023, Zhang et al., 2023, Sypetkowski et al., 2024, Shoghi et al., 2024) and material design (Batatia et al., 2024), GNNs can be applied to the shared feature space of atoms, acting as GFMs. For non-featurized graphs, GNNs equipped with labeling tricks (Zhang et al., 2021) have shown promise as generalizable link predictors in homogeneous graphs (Dong et al., 2024) and multi-relational knowledge graphs (Galkin et al., 2024). Another line of research (Zhao et al., 2023, Tang et al., 2023, Chai et al., 2023, Fatemi et al., 2024, Perozzi et al., 2024, Liu et al., 2024) focuses on text-attributed graphs (Yan et al., 2023), where features are represented as, or converted into, text. Here, large language models (LLMs) serve as universal featurizers by verbalizing the (sub)graph structure and the associated task in natural language. However, it is unclear whether the sequential nature of LLMs is suitable for graphs with permutation invariance.
In summary, while existing GFMs have shown promising results, their generalization capabilities often rely on strong assumptions over the training and test graphs, particularly the presence of a shared input space. In contrast, GraphAny is the first attempt to the more general and challenging problem of generalizing across arbitrary graphs. This broader scope opens up new possibilities for transferring knowledge between different graph domains, such as from knowledge graphs to e-commerce graphs. Additionally, we believe our proposed model, along with essential generalization properties like permutation invariance and dimensional robustness, provide a solid foundation for future research on powerful and versatile GFMs.
3 GraphAny: Fully-inductive Node Classification on Any Graph
@@
Our goal is to devise a fully-inductive model that can perform inductive inference on any new graph with arbitrary feature and label spaces, typically different from the ones associated with the training graph (Figure 2). Here we propose such a solution GraphAny, which consists of two main components (Figure 3): a LinearGNN and an attention module. Each LinearGNN provides a basic solution to inductive inference on new graphs with arbitrary feature and label spaces, while the attention module learns to combine multiple LinearGNNs based on inductive features that generalize to new graphs.
@@
Formally, in a semi-supervised node classification task, we are given a graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, typically represented as an adjacency matrix ${\bm{A}}$, and node features ${\bm{X}}\in{\mathbb{R}}^{|{\mathcal{V}}|\times d}$, a set of labeled nodes ${\mathcal{V}}_{L}$ and their labels ${\bm{Y}}_{L}\in{\mathbb{R}}^{|{\mathcal{V}}_{L}|\times c}$, where $c$ is the number of unique label classes. The goal of node classification is to predict the labels $\hat{{\bm{Y}}}$ for all the unlabeled nodes ${\mathcal{V}}_{U}={\mathcal{V}}\setminus{\mathcal{V}}_{L}$ in the graph. In the conventional transductive learning setup, this is performed by training a GNN (e.g. GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018)) on the subset of labeled nodes using standard backpropagation requiring multiple gradient steps. Such a GNN assumes the full graph to be given and typically does not generalize to a new graph out of the box without some forms of re-training or fine-tuning. Conversely, a fully-inductive model is expected to predict the labels $\hat{{\bm{Y}}}$ for any graph without expensive gradient steps. Furthermore, when a new graph is provided, it might have different dimensionality $d^{\prime}$ of the features and number of class labels, $c^{\prime}$, which can also appear in an arbitrary permuted order.
3.1 Inductive Inference with LinearGNNs
@@
A key idea of this paper is to use simple GNN models whose parameters can be expressed analytically. Following existing works that simplifies GNN (Wu et al., 2019, Zhu and Koniusz, 2021, Yoo et al., 2023) by removing non-linearity, we leverage graph convolutions to process node features, followed by a linear layer to predict the labels,
@@
where ${\bm{F}}={\bm{A}}^{k}{\bm{X}}\in{\mathbb{R}}^{|{\mathcal{V}}|\times d}$ is the processed features and ${\bm{W}}\in{\mathbb{R}}^{d\times c}$ is the weight of the linear layer. Originally, the parameters of most existing GNNs are trained to minimize a cross-entropy loss on node classification, which does not have analytical solution and requires gradient descent to learn the weights. Alternatively, we propose to use a mean-squared error loss for optimizing the weights:
@@
where we use $\hat{{\bm{Y}}_{L}}$ to denote model predictions on the set of labeled nodes. The benefit of this approximation is that now we have an analytical solution for the optimal weights ${\bm{W}}^{*}$:
@@
We term this architecture LinearGNN, as it approximates the prediction of linear GNNs (like SGC (Wu et al., 2019)) with a single forward pass. Its main advantage is that it requires no training, which makes it significantly more computationally efficient (see Section 3.3). Note that the derivation of LinearGNNs is independent of the graph convolution operation, and can be applied to other linear convolution operations as well. We stress that while we do not expect LinearGNNs to outperform existing transductive models on node classification, they provide a simple basic module for inductive inference. In Section 3.2, we will see how to combine multiple LinearGNNs to create a stronger model with inductive attention that generalizes to new graphs.
3.2 Learning Inductive Attention over LinearGNN Predictions
Although LinearGNNs provide basic solutions to inductive inference on new graphs, the learned weights are still transductive, i.e. specific to the training graph, and do not capture transferable knowledge that can be applied to unseen graphs. Besides, our experiments (Figure 7) suggest that different graphs may require LinearGNNs with convolution operations. Hence, a natural way to incorporate fully-inductive learning is to add an inductive attention module over a set of multiple LinearGNNs. Let $\alpha_{u}^{(1)},\alpha_{u}^{(2)},...,\alpha_{u}^{(t)}$ denote the node-level attention over $t$ LinearGNNs. We generate the final prediction as a combination of all LinearGNN predictions: